OpenStudy (anonymous):

which ordered pair is a solution of " 5x+y=-25" a. 0,-5 b. -6,0 c. 0,-6 d. -5,0

OpenStudy (anonymous):

Plug in the values of x and y into the formula, and if both sides are equal, it is right. a. 5(0)+(-5)=-25 0-5=-25 -5=-25 They are not equal, we can rule out A

OpenStudy (anonymous):

b. 5(-6)+0=-25 -30=-25 They are not equal, we can rule out B

OpenStudy (anonymous):

c. 5(0)-6=-25 0-6=-25 -6=-25 They are not equal, we can rule out C

OpenStudy (anonymous):

d. 5(-5)+0=-25 -25=-25 Equal!! D is the answer

OpenStudy (anonymous):

thank you ! :)

OpenStudy (anonymous):

np